When an observation that is much larger than the rest of the data is added to a data​ set, the value of the mean will?
Answer #1

solution:

answer - Increase.

e.g.

2,3,4,5,6. mean = 20/5 =4; mean =4.

now introduce a number which is much greater than these data ,let's say 1000

now data points are 2,3,4,5,6,1000

so, mean = 1020/6 = 170

mean = 170 .

so, mean Increases.

please give me thumb up